OpenStudy (anonymous):

trying to solve this..... x^2+11x+30/x-14=-1

OpenStudy (anonymous):

Start by multiplying both sides by x-14

OpenStudy (anonymous):

ok and that cancels out on the right side right?

OpenStudy (anonymous):

Cancels out on the left, right side becomes -x+14 so now you have x^2+11x+30=-x+14 now move everything to the left and factor

OpenStudy (anonymous):

ok well i got x^2+12x+16

OpenStudy (anonymous):

Yes, and how does that factor

OpenStudy (anonymous):

idk lol

OpenStudy (anonymous):

At least try lol. It doesnt factor nicely, so we could use the quadratic formula or use completing the square method, which one do you like better

OpenStudy (anonymous):

i tried completing the square and the quadratic formula but i dont know how to do the ones that dont work out evenly lol

OpenStudy (anonymous):

Using completing the square x^2+12x+16=0 x^2+12x+36-36+16=0 (x+6)^2-20=0 (x+6)^2=20 Do you know how to do it from there?